Ticket Number: HEA-NOC/20080415-2 Ticket Status: CLOSE
Ticket Type: scheduled Resolver: HEA-NOC
Ticket Opened: 20080415 09:27 UTC+1 Problem Start: 20080415 08:00 UTC+1
Ticket Closed: 20080416 09:58 UTC+1 Problem End: 20080415 09:00 UTC+1
Migration of CAO's current connections to HEAnet to new IP Network
No service interruption is expected, BGP sessions need to be reset and
will be done gradually. A side effect of this migration is asymmetrical
routing while convergence takes place.
Once the changes are completed, traffic will use the new connections,
continuing to have the current ones as backup. Eventually these will be
Contingency plan is to revert changes to existing setup by disabling any
routing over the new links.
20080414 16:00 GL Confirmed with CAO that the migration will take place on
Tuesday 15th April at 8am. Configurations in place.
20080414 08:30 GL Migration completed. CAO's traffic is now routing via a
CRS router in Citywest with old link to Citywest and Galway acting as
secondary and tertiary respectively and will be decommissioned in the next
Monitoring links for 24 hours.
20080416 09:57 GL Link stable, closing ticket
Time to Fix:
This ticket can be monitored at http://www.hea.net/tickets/20080415-2
HEAnet Limited. Registered in Ireland, No. 275301.